OpenStudy (anonymous):

find the following sqrt(x+9)^2

OpenStudy (anonymous):

x + 9 ?

myininaya (myininaya):

\[\sqrt{(x+9)^2}=|x+9|=\pm (x+9)\]

OpenStudy (anonymous):

|x+9|

myininaya (myininaya):

\[|x+9|=x+9 \text{ when } x+9 \ge 0\] \[|x+9|=-(x+9) \text{ when } x+9<0\]

OpenStudy (anonymous):

is it correct to write \[\pm(x+9)\] we don't know what x is

myininaya (myininaya):

lol

OpenStudy (anonymous):

i dont know what x is either

OpenStudy (anonymous):

if you don't know then answer should be |x+9|

myininaya (myininaya):

x+9 is a line sometimes it is positive and something times it is negative